# Local Intrinsic Dimensionality of Ground Motion Data for Early Detection of Complex Catastrophic Slope Failure

### v1

@@ Abstract.
local intrinsic dimensionality, early anomaly detection, spatiotemporal data analysis, landslide early warning systems
1. Introduction
@@
Local Intrinsic Dimensionality (LID) has emerged as a powerful measure for identifying anomalous or outlying behavior in high-dimensional datasets. By effectively capturing representative dimensions within data, LID has demonstrated significant successes across various scenarios such as intrusion detection for IoT networks (Gorbett et al., 2022), identifying adversarial attacks on neural networks (Ma et al., 2018), and outlier detection for high-dimensional data (Anderberg et al., 2024).
@@
Recent advances in granular media mechanics have also demonstrated the effectiveness of LID in the early detection of impending failures in small scale laboratory tests on granular materials (Zhou et al., 2021; Tordesillas et al., 2022; Leśniewska et al., 2023). At larger scales, such failures manifest as some of the most devastating natural hazards. Among these, catastrophic landslides are becoming increasingly frequent and severe due to climate change, seismic activity, and human intervention (Sim et al., 2022; Ozturk et al., 2022; Haque et al., 2019). They can occur suddenly with little warning, often resulting in significant infrastructure damage, loss of life, and long-term economic disruption (Dai et al., 2002; Lacroix et al., 2020; Sim et al., 2022). This sudden onset underscores the importance of timely and accurate identification of high-risk areas to enable early intervention, save lives, and reduce property loss.
@@
Existing works commonly apply traditional machine learning and statistical techniques, such as $K$-means clustering (Tordesillas et al., 2021) and Empirical Dynamic Quantiles (EDQs) (Peña et al., 2019; Tordesillas et al., 2024), to monitoring data on surface displacement in the precursory failure regime, which has been proved to be one of the most direct indicators for identifying the location and timing of slope collapses (Intrieri et al., 2018; Carlà et al., 2019). However, the high dimensionality and inherent spatiotemporal characteristics of such data often challenge the traditional detection methods which are built upon uni-variate time series analysis. As a result, the original informative, multidimensional data is frequently reduced to single dimensional data, leading to significant information loss and potentially inaccurate detection, in both spatial and temporal aspects. Specifically, clustering approaches primarily focus on the spatial detection of high-risk regions without considering the temporal dependencies inherent in the landslide processes. In contrast, studies on EDQs leverage the temporal dynamics of the displacement data but overlook the spatial correlations inside the data. Therefore, these existing works (1) fail to fully exploit both spatial and temporal dependencies inherited in the spatiotemporal displacement data, (2) fail to identify the sudden changes or accelerations in displacement over time. Subsequently, these limitations can lead to inaccurate detection and delay warnings, missing opportunities for timely intervention.
Figure 1. Detection of location and order of successive failures in distinct sites of a slope. Top: The first failure event $C1$ (blue frame) occurred at $t=3385$. Bottom: Four methods detect impending failure locations at an earlier time step $t=3337$ (2 hours before collapse $C1$). Each monitoring point is colored based on the likelihood of failure at its location from 0 to 1. K-means produce binary results, 0 or 1. EDQs highlight points representative of the dynamics against the lower bound 0 for background. $s$-LID correctly identifies both areas of potential collapse but cannot distinguish their order. Only the proposed $st$-LID can detect and isolate the location of the first failure event $C1$ that is relevant at this time step.
@@
To address the challenge in learning with spatiotemporal data, we extend earlier studies of LID in (Zhou et al., 2021) to establish a LID-based framework for dynamic detection of high-risk areas most prone to slope failure at large field scales, by accounting for the outlyingness of kinematic behavior across multiple spatial and time scales. We first add the velocity ($v_{t}$) at each time stamp into target sample, allowing the algorithm to leverage the variability of the displacement and learn the short-term temporal dependence. Incorporating velocity is crucial, as many landslide early warning frameworks rely on velocity thresholds to trigger alerts and guide response actions (Crosta and Agliardi, 2002; Yu et al., 2025). Unlike displacement, which tracks cumulative movement, velocity reflects the rate of change—making it a more immediate indicator of accelerating instability. Secondly, we propose $t$-LID to measure the outlyingness of a target sample with respect to its $t$ nearest temporal neighbors (historical records), enabling it to capture the long-term dynamics from the entire time series data. Finally, we integrate two LID-based scores to highlight the points that are outlying with respect to both spatial and temporal neighborhoods. Specifically, we apply a sigmoid function to normalize both scores into probabilistic values. Then, we integrate two probabilistic scores via multiplication to emphasize samples where both scores are high. This integration not only reduces false positives but also enables earlier detection of hazardous shifts, providing actionable insights well before visible damage occurs. We name the final score spatiotemporal LID ($st$-LID). Experimental results demonstrate that the proposed $st$-LID can pinpoint the actual landslide location with accuracy and efficiency (up to $22\times$ improvement in precision, up to $8\times$ improvement in lead-time). This highlights the practical value of $st$-LID in real-world landslide early warning systems, with the potential to save lives and protect critical infrastructures.
@@
2. Background & Preliminaries
Although catastrophic landslides are driven by multiple factors, surface displacement remains a reliable and measurable indicator of impending failure (Intrieri et al., 2018; Ma and Mei, 2025; Carlà et al., 2019; Finnegan and Saffer, 2024). In the lead-up to collapse, ground motion exhibits a distinct spatiotemporal clustering dynamics, with high-risk zones increasingly differentiating themselves from more stable areas of the slope (Tordesillas et al., 2021; Zhou et al., 2022). Figure 2 shows a typical example of the displacement time series corresponding to a catastrophic slope collapse. Notably, the displacement of points in the failure region manifest three stages of landslide creep (initial, steady-state and accelerated). These patterns highlight the potential for using machine learning and data mining techniques to enable the automatic early detection of slope failure based on surface displacement data (e.g., (Ma and Mei, 2025; Tordesillas et al., 2024; Carlà et al., 2019)).
Figure 2. Example of a real-world slope Failure (Top) and the displacement time series of each monitored point (Bottom). The failure (a collapse advised by domain expert) is located inside the blue frame. Each time series at the bottom represents the displacements of a monitored point along with the time. Two monitored points (inside and outside the failure region) are marked via $\times$’ with different colors in the site map (their corresponding time series are highlighted with the same colors). The red dashed line indicates the time of failure (ToF). And The purple dashed lines indicates three stages of landslide creep.
2.1. Statistical Solutions
Traditional approaches employ statistical methods like $K$-means (with $K=2$) to divide the monitored area into stable and high-risk regions based on their displacement values. Then, metrics like silhouette score and normalized mutual information (NMI) are applied to measure the clustering quality and consistency, subsequently, confirming the detection of the failure regions (Tordesillas et al., 2021). While $K$-means can generally highlight the high-risk regions based on the clustering outputs, its results are inherently binary and depend heavily on the class distinguishability to produce clear and stable boundaries for failure region. Additionally, without leveraging the temporal dynamics, the clustering results are solely based on displacement values at specific time step. These lead to noisy and discrete detection of failure regions, making it hard to highlight finer-scale regions of potential failure.
@@ 2.1. Statistical Solutions
Where $\mathbb{C}_{m}=\{{\bm{x}}_{t}|1\leq t\leq T,{\bm{x}}_{t}\in\mathbb{R}^{m}\}$ is the candidate set, $\bm{1}$} denotes an $m$-dimensional unit vector, ${\bm{y}}_{t}$ is the observation which candidate $x_{t}$ is tested against to, and $\rho_{q}(a)=q\cdot a^{+}+(1-q)\cdot a^{-}$ measures the asymmetrical deviation with $a^{+}=\max(a,0),a^{-}=\max(-a,0)$ (Peña et al., 2019).
When selecting the top quantile levels (e.g., $q\geq 0.75$) EDQs will identify time series that are consistently reflect high percentile behaviors. Consequently, the potential failure regions can be pinpointed by these time series and their corresponding monitored points (Tordesillas et al., 2024). However, since EDQs only focuses on displacement at each monitored points independently without leveraging the spatial relationship between them, it can be easily biased by inherent noise of the displacement data, leading to sparsely distributed results in the monitored site which are infeasible for further analysis.
2.2. Local Intrinsic Dimensionality
While the traditional approaches directly analyzing the raw data values, recent studies that exploit the latent properties of data has proven to be more effective and efficient for machine learning and data science research (Ma et al., 2018; Chen et al., 2020; Huang et al., 2024). The Intrinsic Dimensionality (ID) is one of the latent characteristics of data. It describes the minimal number of latent factors for data representation without losing significant information.
@@ 2.2. Local Intrinsic Dimensionality
Let $\mathcal{F}(r)$ be a real-valued, non-zero function around an open interval of $r\in\mathbb{R}$. If $\mathcal{F}$ is continuously differentiable at $r$, then the LID of $\mathcal{F}$ at $r$ can be defined (Houle, 2017):
(1) $\text{LID}_{\mathcal{F}(r)}:=\lim_{\Delta r\rightarrow 0^{+}}\frac{\ln(\mathcal{F}(r+\Delta r))}{\ln(1+\frac{\Delta r}{r})}=r\frac{\mathcal{F}^{\prime}(r)}{\mathcal{F}(r)},$
@@
Consequently, the LID of a given queried sample $x$ can be derived from (1), letting $r>0$ be a continuous random variable representing the distance from $x$ to its neighbors (other data samples), and $\mathcal{F}(r)$ be the cumulative distribution of $r$ (Zhou et al., 2021):
(2) $\text{LID}(x)=\lim_{r\rightarrow 0}\text{LID}_{\mathcal{F}(r)}.$
Several estimators are proposed to approximate the unknown function $\mathcal{F}$, making LID computable. Among these estimators, Maximum Likelihood Estimator (MLE) has been broadly applied due to its simplicity and efficiency. This also leads to the development of $s$-LID for failure detection in granular mechanics (Zhou et al., 2021): given a data sample at monitored point $p$ with the displacement value of $x_{t}^{p}$, MLE estimates the LID of ${\bm{x}}_{t}=\langle x_{t}^{p}\rangle$ at a fixed time step $t$ based on its distance to its $s$ nearest neighbors in kinematic space (absolute difference of displacement values):
(3) $s\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{s}\sum_{i=1}^{s}\frac{dist_{i}({\bm{x}}_{t})}{dist_{s}({\bm{x}}_{t})}\right)^{-1},$
@@ 3. Methodology
3.1. Leveraging Velocity in LID Calculation
@@
Figure 3. Example high-risk areas and corresponding time series with & without velocity in $s$-LID computation. Top: Site map at last time step ($t=4983$), with two monitored points highlighted by top $s$-LID values, representing the high-risk locations. $\star$ is the $s$-LID without velocity, $\star$ is the $s$-LID with velocity. Bottom: Displacement time series for each monitored point, with the same color.
@@
3.2. Spatial Fusion of LID
Although $s$-LID measures the query point’s characteristic against its spatial neighbors, it finds the nearest neighbors in the kinematic space, without involving the spatial dependencies in physical space. To leverage the spatial relationships of all monitored points during the LID calculation, inspired by Bayesian LID algorithm (Joukhadar et al., 2024), we integrate the $s$-LID results in a small neighborhood of the query point, enhancing the LID score with actual spatial dynamics. Figure 4 depicts the underlying idea, where the query point (red) is updated based on its neighbors (green). Note that the actual number of neighbors is decided during the experiments.
Figure 4. Spatial fusion of $s$-LID. Red dot in the middle is the query point and its neighbors are colored with green. The final LID value of the query point is calculated based on its own observation and the pre-calculated $s$-LID of these neighbors at last time step.
@@ 3.2. Spatial Fusion of LID
(5) $f^{pool}_{t}(l|{\bm{x}}_{t}):=\left(\int_{0}^{\infty}\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}dl\right)\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}.$
@@
(6) $f^{pool}=Ga(\alpha_{p}=\sum_{i=1}^{k}w_{i}\alpha_{i},\beta_{p}=\sum_{i=1}^{k}w_{i}\beta_{i}).$
@@
(7) $\alpha_{p}$ $=\frac{\mu^{2}}{\sigma^{2}},\quad\beta_{p}=\frac{\mu}{\sigma^{2}}.$
@@
(8) $\alpha_{o}=k,\beta_{o}=-\sum_{i=1}^{k}\frac{dist_{i}({\bm{x}}_{t})}{dist_{k}({\bm{x}}_{t})}.$
@@
3.3. Temporally Informed LID
Thirdly, to learn the long-term dynamics of the time series data, we propose a temporal version of LID namely $t$-LID. Same as $s$-LID, $t$-LID is also estimated based on the distance expansion rate. Rather than spatial neighborhood, $t$-LID focuses on temporal neighborhood of the target sample, on another word, the historical values of the given sample until the analyzing time step $t$. Notice that, the raw displacement values are highly non-stationary across the temporal dimension (see Figure 2, 3). To mitigate the outlyingness contributed by such non-stationarity, we compute $t$-LID solely on the velocity values. This method also aligns with the common practice for de-trending non-stationary time series (Mills, 2019). Hence, given a time series of samples at a fixed monitored point $l$: ${\bm{x}}_{1:t}=\{{\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{t}\},{\bm{x}}_{t}=\langle v_{t}^{l}\rangle$, we can write $t$-LID of the queried sample at current time step ${\bm{x}}_{t}$ as:
(9) $t\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{t}\sum_{i=1}^{t}\frac{dist_{i}({\bm{x}}_{t})}{dist_{t}({\bm{x}}_{t})}\right)^{-1},$

### v2

@@ Abstract
1 Introduction
@@
Local Intrinsic Dimensionality (LID) has emerged as a powerful measure for identifying anomalous or outlying behavior in high-dimensional datasets. By effectively capturing representative dimensions within data, LID has demonstrated significant successes across various scenarios such as intrusion detection for IoT networks [7], identifying adversarial attacks on neural networks [15], and outlier detection for high-dimensional data [1].
@@
Recent advances in granular media mechanics have also demonstrated the effectiveness of LID in the early detection of impending failures in small scale laboratory tests on granular materials [26, 23, 14]. At larger scales, such failures manifest as some of the most devastating natural hazards. Among these, catastrophic landslides are becoming increasingly frequent and severe due to climate change, seismic activity, and human intervention [20, 18, 8]. They can occur suddenly with little warning, often resulting in significant infrastructure damage, loss of life, and long-term economic disruption [5, 13, 20]. This sudden onset underscores the importance of timely and accurate identification of high-risk areas to enable early intervention, save lives, and reduce property loss.
@@
Existing works commonly apply traditional machine learning and statistical techniques, such as $K$-means clustering [21] and Empirical Dynamic Quantiles (EDQs) [19, 22], to monitoring data on surface displacement in the precursory failure regime, which has been proved to be one of the most direct indicators for identifying the location and timing of slope collapses [11, 2]. However, the high dimensionality and inherent spatiotemporal characteristics of such data often challenge the traditional detection methods which are built upon uni-variate time series analysis. As a result, the original informative, multidimensional data is frequently reduced to single dimensional data, leading to significant information loss and potentially inaccurate detection, in both spatial and temporal aspects. Specifically, clustering approaches primarily focus on the spatial detection of high-risk regions without considering the temporal dependencies inherent in the landslide processes. In contrast, studies on EDQs leverage the temporal dynamics of the displacement data but overlook the spatial correlations inside the data. Therefore, these existing works (1) fail to fully exploit both spatial and temporal dependencies inherited in the spatiotemporal displacement data, (2) fail to identify the sudden changes or accelerations in displacement over time. Subsequently, these limitations can lead to inaccurate detection and delay warnings, missing opportunities for timely intervention.
@@
To address the challenge in learning with spatiotemporal data, we extend earlier studies of LID in [26] to establish a LID-based framework for dynamic detection of high-risk areas most prone to slope failure at large field scales, by accounting for the outlyingness of kinematic behavior across multiple spatial and time scales. We first add the velocity ($v_{t}$) at each time stamp into target sample, allowing the algorithm to leverage the variability of the displacement and learn the short-term temporal dependence. Incorporating velocity is crucial, as many landslide early warning frameworks rely on velocity thresholds to trigger alerts and guide response actions [4, 24]. Unlike displacement, which tracks cumulative movement, velocity reflects the rate of change—making it a more immediate indicator of accelerating instability.
Figure 1: Detection of location and order of successive failures in distinct sites of a slope. Top: The first failure event $C1$ (blue frame) occurred at $t=3385$. Bottom: Four methods detect impending failure locations at an earlier time step $t=3337$ (2 hours before collapse $C1$). Each monitoring point is colored based on the likelihood of failure at its location from 0 to 1. K-means produce binary results, 0 or 1. EDQs highlight points representative of the dynamics against the lower bound 0 for background. $s$-LID correctly identifies both areas of potential collapse but cannot distinguish their order. Only the proposed $st$-LID can detect and isolate the location of the first failure event $C1$ that is relevant at this time step.
Secondly, we propose $t$-LID to measure the outlyingness of a target sample with respect to its $t$ nearest temporal neighbors (historical records), enabling it to capture the long-term dynamics from the entire time series data. Finally, we integrate two LID-based scores to highlight the points that are outlying with respect to both spatial and temporal neighborhoods. Specifically, we apply a sigmoid function to normalize both scores into probabilistic values. Then, we integrate two probabilistic scores via multiplication to emphasize samples where both scores are high. This integration not only reduces false positives but also enables earlier detection of hazardous shifts, providing actionable insights well before visible damage occurs. We name the final score spatiotemporal LID ($st$-LID). Experimental results demonstrate that the proposed $st$-LID can pinpoint the actual landslide location with accuracy and efficiency (up to $22\times$ improvement in precision, up to $8\times$ improvement in lead-time). This highlights the practical value of $st$-LID in real-world landslide early warning systems, with the potential to save lives and protect critical infrastructures.
@@
2 Background & Preliminaries
Although catastrophic landslides are driven by multiple factors, surface displacement remains a reliable and measurable indicator of impending failure [11, 16, 2, 6]. In the lead-up to collapse, ground motion exhibits a distinct spatiotemporal clustering dynamics, with high-risk zones increasingly differentiating themselves from more stable areas of the slope [21, 25]. Figure 2 shows a typical example of the displacement time series corresponding to a catastrophic slope collapse. Notably, the displacement of points in the failure region manifest three stages of landslide creep (initial, steady-state and accelerated). These patterns highlight the potential for using machine learning and data mining techniques to enable the automatic early detection of slope failure based on surface displacement data (e.g., [16, 22, 2]).
Figure 2: Example of a real-world slope Failure (Top) and the displacement time series of each monitored point (Bottom). The failure (a collapse advised by domain expert) is located inside the blue frame. Each time series at the bottom represents the displacements of a monitored point along with the time. Two monitored points (inside and outside the failure region) are marked via $\times$’ with different colors in the site map (their corresponding time series are highlighted with the same colors). The red dashed line indicates the time of failure (ToF). And The purple dashed lines indicates three stages of landslide creep.
2.1 Statistical Solutions
Traditional approaches employ statistical methods like $K$-means (with $K=2$) to divide the monitored area into stable and high-risk regions based on their displacement values. Then, metrics like silhouette score and normalized mutual information (NMI) are applied to measure the clustering quality and consistency, subsequently, confirming the detection of the failure regions [21]. While $K$-means can generally highlight the high-risk regions based on the clustering outputs, its results are inherently binary and depend heavily on the class distinguishability to produce clear and stable boundaries for failure region. Additionally, without leveraging the temporal dynamics, the clustering results are solely based on displacement values at specific time step. These lead to noisy and discrete detection of failure regions, making it hard to highlight finer-scale regions of potential failure.
@@ 2.1 Statistical Solutions
Where $\mathbb{C}_{m}=\{{\bm{x}}_{t}|1\leq t\leq T,{\bm{x}}_{t}\in\mathbb{R}^{m}\}$ is the candidate set, $\bm{1}$} denotes an $m$-dimensional unit vector, ${\bm{y}}_{t}$ is the observation which candidate $x_{t}$ is tested against to, and $\rho_{q}(a)=q\cdot a^{+}+(1-q)\cdot a^{-}$ measures the asymmetrical deviation with $a^{+}=\max(a,0),a^{-}=\max(-a,0)$ [19].
When selecting the top quantile levels (e.g., $q\geq 0.75$) EDQs will identify time series that are consistently reflect high percentile behaviors. Consequently, the potential failure regions can be pinpointed by these time series and their corresponding monitored points [22]. However, since EDQs only focuses on displacement at each monitored points independently without leveraging the spatial relationship between them, it can be easily biased by inherent noise of the displacement data, leading to sparsely distributed results in the monitored site which are infeasible for further analysis.
2.2 Local Intrinsic Dimensionality
While the traditional approaches directly analyzing the raw data values, recent studies that exploit the latent properties of data has proven to be more effective and efficient for machine learning and data science research [15, 3, 10]. The Intrinsic Dimensionality (ID) is one of the latent characteristics of data. It describes the minimal number of latent factors for data representation without losing significant information.
@@ 2.2 Local Intrinsic Dimensionality
Let $\mathcal{F}(r)$ be a real-valued, non-zero function around an open interval of $r\in\mathbb{R}$. If $\mathcal{F}$ is continuously differentiable at $r$, then the LID of $\mathcal{F}$ at $r$ can be defined [9]:
$\text{LID}_{\mathcal{F}(r)}:=\lim_{\Delta r\rightarrow 0^{+}}\frac{\ln(\mathcal{F}(r+\Delta r))}{\ln(1+\frac{\Delta r}{r})}=r\frac{\mathcal{F}^{\prime}(r)}{\mathcal{F}(r)},$ (1)
@@
Consequently, the LID of a given queried sample $x$ can be derived from (1), letting $r>0$ be a continuous random variable representing the distance from $x$ to its neighbors (other data samples), and $\mathcal{F}(r)$ be the cumulative distribution of $r$ [26]:
$\text{LID}(x)=\lim_{r\rightarrow 0}\text{LID}_{\mathcal{F}(r)}.$ (2)
Several estimators are proposed to approximate the unknown function $\mathcal{F}$, making LID computable. Among these estimators, Maximum Likelihood Estimator (MLE) has been broadly applied due to its simplicity and efficiency. This also leads to the development of $s$-LID for failure detection in granular mechanics [26]: given a data sample at monitored point $p$ with the displacement value of $x_{t}^{p}$, MLE estimates the LID of ${\bm{x}}_{t}=\langle x_{t}^{p}\rangle$ at a fixed time step $t$ based on its distance to its $s$ nearest neighbors in kinematic space (absolute difference of displacement values):
$s\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{s}\sum_{i=1}^{s}\frac{dist_{i}({\bm{x}}_{t})}{dist_{s}({\bm{x}}_{t})}\right)^{-1},$ (3)
@@ 3 Methodology
3.1 Leveraging Velocity in LID Calculation
@@
Figure 3: Example high-risk areas and corresponding time series with & without velocity in $s$-LID computation. Top: Site map at last time step ($t=4983$), with two monitored points highlighted by top $s$-LID values, representing the high-risk locations. $\star$ is the $s$-LID without velocity, $\star$ is the $s$-LID with velocity. Bottom: Displacement time series for each monitored point, with the same color.
@@
3.2 Spatial Fusion of LID
Although $s$-LID measures the query point’s characteristic against its spatial neighbors, it finds the nearest neighbors in the kinematic space, without involving the spatial dependencies in physical space. To leverage the spatial relationships of all monitored points during the LID calculation, inspired by Bayesian LID algorithm [12], we integrate the $s$-LID results in a small neighborhood of the query point, enhancing the LID score with actual spatial dynamics. Figure 4 depicts the underlying idea, where the query point (red) is updated based on its neighbors (green). Note that the actual number of neighbors is decided during the experiments.
Figure 4: Spatial fusion of $s$-LID. Red dot in the middle is the query point and its neighbors are colored with green. The final LID value of the query point is calculated based on its own observation and the pre-calculated $s$-LID of these neighbors at last time step.
@@ 3.2 Spatial Fusion of LID
$f^{pool}_{t}(l|{\bm{x}}_{t}):=\left(\int_{0}^{\infty}\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}dl\right)\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}.$ (5)
@@
$f^{pool}=Ga(\alpha_{p}=\sum_{i=1}^{k}w_{i}\alpha_{i},\beta_{p}=\sum_{i=1}^{k}w_{i}\beta_{i}).$ (6)
@@
$\alpha_{p}$ $=\frac{\mu^{2}}{\sigma^{2}},\quad\beta_{p}=\frac{\mu}{\sigma^{2}}.$ (7)
@@
$\alpha_{o}=k,\beta_{o}=-\sum_{i=1}^{k}\frac{dist_{i}({\bm{x}}_{t})}{dist_{k}({\bm{x}}_{t})}.$ (8)
@@
3.3 Temporally Informed LID
Thirdly, to learn the long-term dynamics of the time series data, we propose a temporal version of LID namely $t$-LID. Same as $s$-LID, $t$-LID is also estimated based on the distance expansion rate. Rather than spatial neighborhood, $t$-LID focuses on temporal neighborhood of the target sample, on another word, the historical values of the given sample until the analyzing time step $t$. Notice that, the raw displacement values are highly non-stationary across the temporal dimension (see Figure 2, 3). To mitigate the outlyingness contributed by such non-stationarity, we compute $t$-LID solely on the velocity values. This method also aligns with the common practice for de-trending non-stationary time series [17]. Hence, given a time series of samples at a fixed monitored point $l$: ${\bm{x}}_{1:t}=\{{\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{t}\},{\bm{x}}_{t}=\langle v_{t}^{l}\rangle$, we can write $t$-LID of the queried sample at current time step ${\bm{x}}_{t}$ as:
$t\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{t}\sum_{i=1}^{t}\frac{dist_{i}({\bm{x}}_{t})}{dist_{t}({\bm{x}}_{t})}\right)^{-1},$ (9)